OpenStudy (johnnydicamillo):

Solve for x - 7/x-6 = 7/2x-12 + 4

OpenStudy (igreen):

Is it \(\dfrac{-7}{x-6}\) or \(\dfrac{-7}{x} - 6\)

OpenStudy (johnnydicamillo):

first one

OpenStudy (igreen):

Okay.

OpenStudy (johnnydicamillo):

\[7/(2x-12) + 4\] for the second part

OpenStudy (igreen):

Okay.

OpenStudy (igreen):

OKay, your answers gonna be \(\dfrac{27}{8}\) or \(3\dfrac{3}{8}\).

OpenStudy (igreen):

*Okay

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'.

OpenStudy (johnnydicamillo):

can you explain the process to me

OpenStudy (igreen):

Sorry, I g2g work on my History exam D: